OpenStudy (anonymous):

Try for fun, Solve for x where (x^3)-15x=126

OpenStudy (anonymous):

Pl dont post the solution. Post result only OR msg me.

OpenStudy (anonymous):

Posting for posterity. Would like to see a hand-done solution.

OpenStudy (amistre64):

id like to see a toe done solution meself :)

OpenStudy (amistre64):

3x^2-15 = 0 when x = +-sqrt(15/3) 6x = 0 when x=0 , so this changes concavity at x=0 we have no undefined points id say it has to be when x is greater than sqrt(15/3) so far :)

OpenStudy (anonymous):

I wonder if you can take the mean distance between change points in concavity, and find the \(x\) that way. That's what I tried, but I have a headache and math seems to be difficult right now. :c

OpenStudy (shubhamsrg):

x^3 - 15x =126 => x^3 - 15 x - 126 =0 =>x^3 + 6x^2 + 21x - 6x^2 - 36x - 126=0 should be easier now..

OpenStudy (shubhamsrg):

another way can be rational root theorem..

OpenStudy (ajprincess):

x=6 x=(+/-sqrt-12)-3=(+/-)(2*sqrt-3)-3

OpenStudy (anonymous):

i got x = 6 by noticing that if you add 90 = 15(6) to 126 you get a lovely cube :)